A student has $8 to spend on a phone call. The cost of a call is $.34 for the first minute and $.24 for each additional minute. How long can the student talk on the phone?
Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

A student has $8 to spend on a phone call. The cost of a call is $.34 for the first minute and $.24 for each additional minute. How long can the student talk on the phone?
--------------
Cost = 34 + 24m = 800
Solve for m

Answer by Theo(13342)

the cost of a call is .34 for the first minute and .24 for each additional minute.
the student has 8.00 to spend.
the cost would be .34 + .24 * (x-1) for x >= 1.
when x = 1, the cost is .34 + .24 * 0 = .34
for x = 2, the cost is .34 + .24 * 1 = .58
for x = 3, the cost is .34 + .24 * 2 = .82
etc.
the equation can be graphed.
it is shown below:

to answer your question, let the equation be y = .34 + .24 * (x-1) for x greater than equal to 1.
when y = 8, the equation becomes 8 = .34 + .24 * (x-1).
subtract .34 from both sides of the equation to get 7.66 = .24 * (x-1)
divide both sides of the equation by .24 to get 31.916666... = x - 1
add 1 to both sides of the equation to get 32.916666.... = x
assuming x can only be an integer and must be less than or equal to 32.916666..., then, when x = 32, you will pay .34 + 31 * .24 = 7.78 for 32 minutes of calls and you will have .22 left over.